OpenStudy (anonymous):

A local hamburger shop sold a combined total of 639 hamburgers and cheeseburgers on Monday. There were 61 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday? need it NOW D:

OpenStudy (anonymous):

suppose x is the number of hamburgers sold suppose y is the number of cheeseburgers what is the relationship between x and y.

OpenStudy (anonymous):

let H = # of hamburgers C = # of cheeseburgers.. C + 61 = H... Now, construct your equation! H + C = 639

OpenStudy (anonymous):

Using substitution: (C + 61) + C = 639 Then, algebra: 2c + 61 = 639... 2C = 578 C = 289....

OpenStudy (anonymous):

If C = 289, then H = 350

OpenStudy (anonymous):

Understood. Thanks :)

OpenStudy (anonymous):

Great! Best of luck...